OpenStudy (anonymous):

(-1, 3) and perpendicular to y = -1/3 x + 7. halp

OpenStudy (anonymous):

i need steps, not just the answer. i have no idea how to do this.

OpenStudy (kymber):

Is it asking you to find the slope?

OpenStudy (anonymous):

no idea, thats all it says. the choices are y=3x-12 y=3x+6 y=3x-6 y=3x